### v1

# M4-SAR: A Multi-Resolution, Multi-Polarization, Multi-Scene, Multi-Source Dataset and Benchmark for Optical-SAR Fusion Object Detection

Abstract
Single-source remote sensing object detection using optical or SAR images struggles in complex environments. Optical images offer rich textural details but are often affected by low-light, cloud-obscured, or low-resolution conditions, reducing the detection performance. SAR images are robust to weather, but suffer from speckle noise and limited semantic expressiveness. Optical and SAR images provide complementary advantages, and fusing them can significantly improve the detection accuracy. However, progress in this field is hindered by the lack of large-scale, standardized datasets. To address these challenges, we propose the first comprehensive dataset for optical-SAR fusion object detection, named Multi-resolution, Multi-polarization, Multi-scene, Multi-source SAR dataset (M4-SAR). It contains 112,184 precisely aligned image pairs and nearly one million labeled instances with arbitrary orientations, spanning six key categories. To enable standardized evaluation, we develop a unified benchmarking toolkit that integrates six state-of-the-art multi-source fusion methods. Furthermore, we propose E2E-OSDet, a novel end-to-end multi-source fusion detection framework that mitigates cross-domain discrepancies and establishes a robust baseline for future studies. Extensive experiments on M4-SAR demonstrate that fusing optical and SAR data can improve $mAP$ by 5.7% over single-source inputs, with particularly significant gains in complex environments. The dataset and code are publicly available at https://github.com/wchao0601/M4-SAR.
1 Introduction
With the rapid development of deep learning techniques, single-source detection (e.g., optical [27; 35] or synthetic aperture radar (SAR) [20; 55]), and multi-source detection (optical-SAR [26; 71]) have garnered significant attention in the remote sensing field. These methods have demonstrated remarkable effectiveness in applications, including disaster monitoring and urban planning. As shown in Figure 1 (c), optical images provide high-resolution visual cues with rich texture, color, and spectral information, facilitating fine-grained feature extraction and semantic interpretation under clear atmospheric conditions. However, their detection performance degrades markedly under low-light, cloud-obscured, or low-resolution conditions (see Figure 1 (a)). Despite these drawbacks, the intuitive appearance and broad availability of optical data have fueled significant advancements in optical image object detection [31; 77; 4; 7]. In contrast, SAR sensors enable all-weather, all-day imaging to cloud cover and illumination changes. However, the inherent speckle noise and low contrast arising from microwave backscatter present significant obstacles to precise localization and classification. Moreover, constructing large-scale SAR datasets [28; 76; 62] is hindered by the high cost of data acquisition and the laborious fine-grained annotation. Although optical and SAR images exhibit strong complementarity in spatial resolution and environmental robustness, existing single-source [49; 74; 36] and multi-source approaches [18; 2; 54] face challenges in addressing cross-domain distribution shifts and suffer from the absence of a standardized evaluation benchmark. To enable robust detection in complex scenarios, there is an urgent need for a large-scale, precisely aligned optical-SAR fusion detection dataset and a standardized benchmarking toolkit.
Currently, a major bottleneck in optical–SAR fusion object detection lies in the lack of large-scale, high-quality, and standardized datasets. As summarized in Table 1, most existing SAR benchmarks are limited to single-source data, suffering from limited scale, narrow category coverage, or inconsistent annotations. For example, HRSID [61], SADD [73], SSDD [75], SIVED [29], and ShipDataset [59] exhibit notable limitations in terms of category diversity and dataset scale. In recent years, the emergence of larger-scale datasets such as SARDet-100K [28], FAIR-CSAR [62], and RSAR [76] has partially addressed these limitations. However, these datasets are often synthesized by standardization of data from multiple sources, potentially introducing bias during data preprocessing and compromising the fairness of model evaluation. Moreover, constructing paired optical–SAR datasets is prohibitively expensive: acquiring SAR data is already costly, and the acquisition of spatio-temporally aligned optical images is even more difficult, further limiting the construction of multi-source datasets. OGSOD-1.0 [55] is currently the only optical-SAR multi-source dataset, but it is designed to serve mainly cross-modal knowledge distillation, and the test set contains only SAR images, thereby limiting its suitability for comprehensive multi-source fusion object detection. To address this issue, we constructed the first optical-SAR fusion object detection dataset (M4-SAR) covering multi-resolution, multi-polarization, multi-scene, and multi-source using publicly available optical and SAR data provided by Sentinel-1 [50] and Sentinel-2 [9] satellites. To overcome SAR annotation challenges, we propose a semi-supervised optical-assisted labeling strategy that utilizes optical images’ semantic richness to substantially improve annotation quality. As illustrated in Figure 1 and Table 1, M4-SAR includes six key target categories with 112,184 image pairs and 981,862 instances. M4-SAR surpasses existing datasets in instance volume and alignment precision, establishing a robust benchmark for multi-source fusion in complex scenarios.
Figure 1: Illustrative examples of scenes and six key target categories in the proposed M4-SAR dataset, accompanied by instance size and aspect ratio distributions.
Another important challenge is the lack of publicly available algorithms tailored for optical–SAR fusion object detection. Most existing studies are based on single-source datasets or evaluated on limited scenarios, making it difficult to conduct fair performance comparisons. To bridge this gap, we introduce Multi-Source Rotated Object Detector (MSRODet), an open-source toolkit that integrates state-of-the-art fusion detectors (e.g., CFT [42], CLANet [18], CSSA [2], CMADet [45], ICAFusion [45], and MMIDet [69]), providing a standardized evaluation framework for researchers utilizing the M4-SAR dataset. MSRODet not only enables fair and reproducible benchmarking across different fusion methods but also lays a solid foundation for the development of future multi-source fusion object detection algorithms. More details of MSRODet are presented in Appendix B.
Moreover, extensive experiments demonstrate that the intrinsic differences in the imaging mechanisms of optical and SAR sensors lead to significant domain gaps (Figure 4 (a)), thereby constraining the performance of optical–SAR fusion object detection. To address this issue, we propose the first end-to-end framework for optical–SAR fusion object detection, termed E2E-OSDet. This framework effectively fuses features from both modalities, alleviates cross-modal domain gaps, and results in significant improvements in both inference speed and detection performance.
This work offers substantial contributions in terms of dataset construction, toolkit development, and algorithm design. The main contributions of this work are summarized as follows:
  • M4-SAR Dataset: M4-SAR is the first optical-SAR fusion object detection dataset with multi-resolution, multi-polarization, and multi-scene characteristics, containing 112,184 image pairs and 981,862 instances. This dataset comprehensively covers complex scenes and diverse target categories frequently encountered in multi-source fusion detection tasks, thereby providing a robust foundation for advancing research in this field.
  • Multi-source Evaluation Toolkit (MSRODet) and E2E-OSDet: We introduce MSRODet, a comprehensive benchmarking toolkit incorporating state-of-the-art fusion methods tailored for optical–SAR object detection, which provides a standardized and reproducible evaluation platform based on the M4-SAR dataset. In addition, we propose E2E-OSDet, a novel end-to-end optical–SAR fusion object detection framework, which effectively mitigates significant cross-domain discrepancies between optical and SAR modalities at the feature level.
Table 1: Comparison between existing optical and SAR object detection datasets and the proposed M4-SAR dataset. SR denotes the spatial resolution of optical images.
Dataset OPT SAR SR & Polarization Image Size Category Instances Images Ins/Img
HRSC2016 [31] $\checkmark$ 0.4M $\sim$ 2M 300 $\sim$ 1,500 1 2,976 1,061 2.80
LEVIR [77] $\checkmark$ 0.2M $\sim$ 1M 600 $\times$ 800 3 11,028 21,952 0.50
DIOR-R [4] $\checkmark$ 0.5M $\sim$ 30M 800 $\times$ 800 20 190,288 23,463 8.11
DOTA-v1.0 [7] $\checkmark$ - 800 $\sim$ 4,000 15 188,282 2,806 67.10
HRSID [61] $\checkmark$ HH, HV, VH, VV 800 $\times$ 800 1 16,969 5,604 3.03
SADD [73] $\checkmark$ HH 224 $\times$ 224 1 7,835 883 8.87
SSDD [75] $\checkmark$ HH, HV, VH, VV 512 $\times$ 512 1 2,587 1,160 2.23
SIVED [29] $\checkmark$ VV, VH 512 $\times$ 512 1 12,013 1,044 11.51
ShipDataset [59] $\checkmark$ HH, HV, VH, VV 256 $\times$ 256 1 50,885 39,729 1.28
MSAR [63] $\checkmark$ HH, HV, VH, VV 256 $\times$ 256 4 65,202 30,158 2.16
SARDet-100K [28] $\checkmark$ HH, HV, VH, VV 512 $\times$ 512 6 245,653 116,598 2.11
FAIR-CSAR [62] $\checkmark$ HH, HV, VH, VV 1,024 $\times$ 1,024 5 349,002 106,672 2.11
RSAR [76] $\checkmark$ HH, HV, VH, VV 512 $\times$ 512 6 183,534 95,842 1.91
OGSOD-1.0 [55] $\checkmark$ $\checkmark$ 20M, VV, VH 256 $\times$ 256 3 48,589 18,331 2.65
M4-SAR (Ours) $\checkmark$ $\checkmark$ 10M, 60M, VV, VH 512 $\times$ 512 6 981,862 112,184 8.75
2 M4-SAR Dataset
Overview of the Proposed Dataset Construction Pipeline. Figure 2 illustrates the overall pipeline of the proposed dataset construction process. In the first stage, we acquire optical and SAR images from Sentinel-1 and Sentinel-2 satellites provided by the European Space Agency (ESA). In the second stage, the acquired images are pre-processed and aligned. After completing the alignment, we crop aligned image pairs into fixed-size patches to generate initial unlabeled samples. In the third stage, we introduce a semi-supervised annotation strategy guided by optical images to ensure high-quality labeling of the image pairs. More dataset details are presented in Appendix A.
Data Collection Procedure. First, we collected the required data via the open-access platform provided by the European Space Agency (ESA). A map-based selection tool was used to specify areas of interest, covering representative coastal cities, including Tianjin, Los Angeles, Hong Kong, and London. We then retrieved SAR images along with their corresponding digital elevation model (DEM) data according to specific criteria. For optical imagery, Sentinel-2 images covering the same regions as the Sentinel-1 data were acquired, specifically including the red (R), green (G), and blue (B) spectral bands. To minimize variations caused by temporal discrepancies between the two modalities, we selected image pairs with a time interval of less than 10 days. This setting ensures that targets of interest (e.g., bridges, oil tanks) maintain structural consistency within the short interval, thereby enhancing the accuracy of subsequent instance-level alignment. The product type for optical images is L2A, with cloud coverage limited to 0–10% to guarantee high visual fidelity. The SAR images are provided in the Single Look Complex (SLC) format to preserve full coherence information. The collected data span the period from 2020 to 2022 and encompass multiple spatial resolutions and polarization modes. The optical images have 10m, 20m, and 60m resolutions, and the SAR data cover both vertical-horizontal (VH) and vertical-vertical (VV) polarizations.
Data Processing. We combined the red, green, and blue channels to generate standard optical images using data processing software. The optical images were provided in three spatial resolutions: 10m, 20m, and 60m. To highlight the complementary characteristics of optical and SAR images, we retained only the high-resolution (10m) and low-resolution (60m) variants. For the SAR data, we first applied multi-look averaging to the SLC images in both the azimuth and range directions to reduce speckle effects and improve radiometric resolution. Subsequently, filtering techniques were employed to suppress speckle noise. Using the corresponding DEM data, we geocoded and radiometrically calibrated the SAR images and conducted geometric alignment and correction to improve spatial localization precision. After preprocessing, the optical and SAR images were aligned based on their geographic coordinates. As the overlap between the two modalities was partial, the aligned image pairs were cropped into 1,024$\times$1,024 patches, and only those with valid overlapping regions were retained for subsequent annotation. Notably, this initial cropping was intended to facilitate efficient manual annotation. After annotation was completed, we further partitioned the annotated images into non-overlapping 512$\times$512 tiles to meet the demands of different downstream applications.
Figure 2: Pipeline for M4-SAR dataset construction.
Data Annotation and Label Mapping. As the acquisition times of Sentinel-1 [50] and Sentinel-2 [9] are not perfectly synchronized, instance-level alignment of dynamic targets (e.g., ships) becomes difficult, which compromises the accuracy of label transfer. To address this issue, we focus on static and semantically meaningful targets, such as bridges, harbors, oil tanks, playgrounds, airports, and wind turbines, and annotate them using oriented bounding boxes. SAR’s unique imaging characteristics cause ambiguous target boundaries, challenging accurate annotation by non-experts. In contrast, optical images provide clearer visual cues that facilitate both target identification and annotation transfer. With precise alignment, annotations derived from optical images can be reliably transferred on SAR images, thereby significantly reducing manual annotation costs.
Given the scale of the dataset, fully manual labeling is both time-consuming and labor-intensive. To address this challenge, we propose a semi-supervised annotation strategy, as illustrated in Figure 2(c). The procedure involves the following steps: First, 5% of the entire dataset is manually annotated, ensuring that each selected image contains all six target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. This subset establishes a high-quality foundation for subsequent annotation. Once manually labeled, a generic detector [21] is trained using this subset (with the same data used for both training and evaluation during the initial phase). The trained model is then employed to generate pseudo-labels for an additional 5% of the data, which are subsequently refined through manual correction and added to the labeled set. This updated set is used to retrain the detector in the next iteration. The annotation process is repeated iteratively until the full dataset is labeled. During this procedure, images without valid target instances are excluded to enhance annotation efficiency and the overall utility of the dataset.
Based on instance-level alignment, we achieve high-quality label transfer, enabling accurate projection of annotation information from optical images to their corresponding SAR counterparts. As a result, we construct M4-SAR, the first large-scale dataset for optical-SAR fusion object detection, which offers a solid foundation for subsequent method evaluation and algorithm development.
Statistical Analysis of Different Modalities. As shown in Table 1, our proposed M4-SAR dataset comprises 112,184 image pairs, 981,862 instances, and six representative target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. Each image has a resolution of 512$\times$512 pixels and contains an average of 8.75 instances, highlighting the dense object distribution typical in remote sensing scenes. As illustrated in Figure 1 (c), rotated bounding boxes offer significantly higher localization accuracy compared to horizontal bounding boxes.
Figure 3: Statistical visualization of category attributes in the proposed M4-SAR dataset. (a) Angle distribution of instances in each category. (b) Percentage of instances per category. (c) Average pixel area of instances per category.
Statistical analysis of category-specific attributes in the M4-SAR dataset is presented as follows:
  • Aspect Ratio: The aspect ratio of an object plays a critical role in detection accuracy. Figure 1 (b) illustrates the distribution of aspect ratios across categories. When the aspect ratio approaches 1, targets tend to appear square, potentially leading to recognition ambiguities. Conversely, extremely large aspect ratios increase the model’s sensitivity to object shapes, thus posing additional challenges to accurate detection.
  • Angle: All target orientation angles are normalized to $\theta$ ($\theta\in[0,\pi/2)$), and their distributions are visualized in Figure 3(a). Notably, offshore wind turbines tend to exhibit square-like appearances, with orientation angles predominantly distributed around 0 and 90 due to the absence of surrounding reference structures and the top-down perspective of satellite imagery. Although oil tanks are circular, their labeled angles are contextually defined based on surrounding spatial features. This highlights the importance of contextual cues in remote sensing images and reflects the task’s sensitivity and complexity.
  • Number of Instances: Figure 3 (b) presents the instance distribution by category, where bridges and oil tanks constitute the majority of instances.
  • Average Area of Objects: Figure 3 (c) shows the average pixel area per category, revealing the scale variability among different object categories in the dataset.
3 Benchmarks
Optical-SAR Fusion Detection Tools (MSRODet). Due to the lack of publicly available methods specifically designed for optical-SAR fusion object detection, conducting a fair performance comparison remains challenging. To address this limitation, we propose an optical-SAR fusion object detection evaluation toolbox (MSRODet) built upon existing multi-source detection frameworks (e.g., RGB-infrared, RGB-depth). MSRODet integrates several representative methods, including CFT [42], CLANet [18], CSSA [2], CMADet [45], ICAFusion [45], and MMIDet [69], thereby providing a standardized and accessible benchmark for evaluating optical-SAR fusion object detection approaches. As shown in Figure 4 (a), empirical analysis reveals that significant domain gaps remain a primary obstacle to effective optical-SAR fusion. These discrepancies primarily arise from fundamental differences in visual characteristics and imaging mechanisms between optical and SAR modalities. Consequently, designing detection frameworks specifically tailored to optical-SAR fusion is essential for enhancing overall detection performance.
Figure 4: (a) Optical-SAR domain gap. (b) Overall framework of the proposed E2E-OSDet.
Figure 5: Architectural details of the proposed FAM, CMIM, and AFM modules.
End-to-End Optical-SAR Fusion Object Detection (E2E-OSDet). Multi-source fusion object detection remains underexplored in the context of optical-SAR fusion, and significant limitations persist in addressing domain discrepancies between optical and SAR data. As shown in Figure 4 (b), to address these challenges, we propose a novel end-to-end optical-SAR fusion object detection framework, termed E2E-OSDet, which systematically addresses cross-modal challenges across three levels: data input, domain alignment, and feature fusion.
Filter Augment Module (FAM): First, to reduce the domain gap between optical and SAR images, we propose a Filter Augment Module (FAM) leveraging classical image filtering operators (e.g., HOG [5], Canny [1], Haar [52], Grad [24], and WST [40]). This module transforms the low-dimensional and sparse SAR representations into a more discriminative high-dimensional space, effectively mitigating the impact of domain differences.
Cross-modal Mamba Interaction Module (CMIM): Second, most existing methods focus on spatial-level fusion but often neglect the discrepancies in the feature domain. Mamba [14] possesses strong state-space modeling capabilities, which enables effective cross-modal sequence-level interactions and alleviates feature domain bias. However, current Mamba-based methods [70; 8] are hindered by complex architectures and intricate interaction mechanisms, limiting their practical applicability. To address this, we propose a Cross-modal Mamba Interaction Module (CMIM), which enables deep cross-modal interactions through simple and efficient operations.
Area-Attention Fusion Module (AFM): Finally, to further enhance the discriminability of fused features across different spatial regions, we introduce an Area-Attention Fusion Module (AFM). This module captures salient features of local targets and strengthens responses in critical regions, thereby enhancing the effectiveness of feature fusion. The detailed structure of each module is illustrated in Figure 5. For additional E2E-OSDet details, please refer to Appendix C.
4 Experiments
4.1 Setups and Implementation Details
Experiments were conducted on two NVIDIA RTX 4090 GPUs. All fusion models were trained over 300 epochs with the SGD optimizer, using a learning rate of 0.01, momentum of 0.937, and weight decay of 5$\times$10-4. The batch size was set to 64, and the input images were resized to a resolution of 512$\times$512. To ensure fair performance evaluation, all methods were evaluated using a unified framework comprising the YOLOv11 [21] backbone and the oriented bounding box (OBB) detection head from YOLOv8 [21]. All fusion models were trained from scratch, without leveraging any pre-trained weights. During evaluation, we adopted the standard COCO evaluation protocol to measure Average Precision (AP), including $AP_{50}$, $AP_{75}$, and mean AP ($mAP$), where $mAP$ denotes the mean AP across IoU thresholds ranging from 0.5 to 0.95. For more details, see Appendix D.
4.2 Quantitative Evaluation
Main Results. We compared multi-source fusion methods (e.g., CFT, CLANet, CSSA, CMADet, ICAFusion, MMIDet, and E2E-MFD) with single-source methods (e.g., RTMDet, PSC, KFIoU, LSKNet and YOLO series) that utilize either optical or SAR data exclusively. As shown in Table 2, the detection accuracy achieved using SAR data alone is considerably lower than that of the optical-only counterpart. This performance gap is primarily attributed to the unique imaging principles of SAR, which hinder the accurate estimation of bounding box shapes—even when targets are visually identifiable. All multi-source fusion methods outperform their corresponding unimodal baselines, demonstrating the effectiveness of cross-modal fusion in leveraging SAR data to complement optical information (and vice versa), while significantly reducing false positives. Notably, our proposed method, E2E-OSDet, achieves the highest performance among all fusion-based approaches, reaching an $mAP$ of 61.4%, which underscores the importance of mitigating domain discrepancies to improve detection accuracy. Furthermore, E2E-OSDet is also highly efficient, with only 27.5M parameters and an inference time of 20.9ms. Finally, we evaluate the generalizability of E2E-OSDet across various detection frameworks, with the results presented in Figure 6(e).
Table 2: Quantitative results on the M4-SAR. O and S denote optical and SAR data, respectively. Inf.T indicates inference time per image. The best two results are highlighted in red and green.
[…]
R-FCOS [49] ICCV’19 $\checkmark$ 31.9 17.6ms 32.8 79.8 29.7 67.0 78.9 54.8 57.2 26.4 30.2
R-ATSS [74] CVPR’20 $\checkmark$ 36.0 19.1ms 33.2 80.0 29.3 66.6 78.3 54.8 57.0 27.1 30.0
YOLOv5 [21] arXiv’20 $\checkmark$ 8.10 7.20ms 61.4 88.0 56.6 76.4 91.3 91.8 77.6 59.8 52.9
Roi-Trans [39] arXiv’21 $\checkmark$ 273 49.3ms 17.2 80.7 9.1 55.2 80.5 25.6 44.2 27.6 27.1
O-RepPoint [25] CVPR’22 $\checkmark$ 36.6 16.1ms 63.3 90.2 55.8 81.7 88.2 90.1 77.1 58.3 53.9
RTMDet [38] arXiv’22 $\checkmark$ 8.86 11.1ms 34.0 79.8 30.3 61.1 79.1 55.9 56.7 26.5 29.2
YOLOv8 [21] arXiv’23 $\checkmark$ 10.0 6.70ms 63.6 88.6 57.9 80.0 91.3 92.5 79.0 61.6 54.4
PSC [66] CVPR’23 $\checkmark$ 31.9 18.3ms 30.7 79.7 25.9 57.4 78.1 49.4 53.5 26.2 28.5
KFIoU [65] ICLR’23 $\checkmark$ - 66.2ms 16.9 79.4 9.1 50.9 80.4 25.2 43.6 19.4 22.4
YOLOv6 [22] ICLR’24 $\checkmark$ 16.0 6.10ms 57.6 82.1 52.7 67.6 89.4 90.4 73.3 64.9 48.8
YOLOv9 [56] ECCV’24 $\checkmark$ 6.40 14.4ms 61.9 87.7 55.3 74.4 90.0 92.4 76.9 58.9 52.5
YOLOv10 [53] NeurIPS’24 $\checkmark$ 7.50 7.80ms 65.2 89.9 58.8 82.3 90.8 93.5 80.1 62.8 55.7
YOLOv11 [21] arXiv’24 $\checkmark$ 9.74 6.50ms 64.3 90.0 57.8 80.7 90.9 91.5 79.2 60.5 54.3
LSKNet [27] IJCV’24 $\checkmark$ 21.8 20.4ms 63.3 90.2 55.8 81.7 88.2 90.1 77.1 58.3 53.9
R-FCOS [49] ICCV’19 $\checkmark$ 31.9 17.6ms 23.9 80.1 27.9 60.1 75.8 43.5 51.9 26.4 28.3
R-ATSS [74] CVPR’20 $\checkmark$ 36.0 19.1ms 25.3 80.3 26.8 62.3 77.0 43.9 52.6 24.9 27.8
YOLOv5 [21] arXiv’20 $\checkmark$ 8.10 7.20ms 40.5 85.3 41.0 64.3 85.5 89.9 67.7 44.0 41.3
Roi-Trans [39] arXiv’21 $\checkmark$ 273 49.3ms 9.1 80.8 9.1 35.6 79.0 18.0 38.6 25.2 24.3
O-RepPoint [25] CVPR’22 $\checkmark$ 36.6 16.1ms 26.2 78.6 21.9 53.7 59.9 48.8 48.2 13.9 20.4
RTMDet [38] arXiv’22 $\checkmark$ 8.86 11.1ms 16.2 71.0 21.1 39.3 68.9 39.6 42.7 20.9 22.3
YOLOv8 [21] arXiv’23 $\checkmark$ 10.0 6.70ms 44.9 85.8 42.5 67.9 87.5 89.2 69.6 46.1 43.2
PSC [66] CVPR’23 $\checkmark$ 31.9 18.3ms 22.3 80.0 20.6 51.9 75.7 39.4 48.3 24.5 26.5
KFIoU [65] ICLR’23 $\checkmark$ - 66.2ms 9.1 72.1 9.1 35.3 78.0 22.7 37.7 17.4 19.8
YOLOv6 [22] ICLR’24 $\checkmark$ 16.0 6.10ms 36.5 79.4 32.8 53.5 85.7 86.7 62.4 37.6 36.5
YOLOv9 [56] ECCV’24 $\checkmark$ 6.40 14.4ms 43.9 84.8 38.4 61.5 87.9 88.8 67.6 44.2 41.4
YOLOv10 [53] NeurIPS’24 $\checkmark$ 7.50 7.80ms 49.8 87.0 45.5 73.4 87.9 90.4 72.3 50.8 46.2
YOLOv11 [21] arXiv’24 $\checkmark$ 9.74 6.50ms 46.9 87.2 44.6 71.1 88.8 89.2 71.3 48.6 44.7
LSKNet [27] IJCV’24 $\checkmark$ 21.8 20.4ms 33.4 80.2 28.4 69.3 76.8 51.2 56.6 27.1 30.4
CFT [42] PR’22 $\checkmark$ $\checkmark$ 53.8 40.6ms 75.8 92.5 61.3 91.6 90.3 96.3 84.6 68.9 59.9
CLANet [18] ACMMM’23 $\checkmark$ $\checkmark$ 48.2 29.1ms 74.8 92.2 60.7 91.3 91.6 97.2 84.6 68.5 59.6
CSSA [2] CVPR’23 $\checkmark$ $\checkmark$ 13.5 12.3ms 73.3 91.7 59.3 88.9 91.6 95.8 83.4 66.4 58.0
CMADet [45] TIV’24 $\checkmark$ $\checkmark$ 41.5 46.7ms 70.9 90.7 52.0 86.4 91.7 97.1 81.5 63.5 55.7
ICAFusion [44] PR’24 $\checkmark$ $\checkmark$ 29.0 23.6ms 74.7 91.9 60.9 91.0 91.8 96.7 84.5 67.3 58.8
MMIDet [69] TCSVT’24 $\checkmark$ $\checkmark$ 53.8 41.9ms 74.9 92.6 61.1 91.7 91.4 97.0 84.8 68.6 59.8
E2E-MFD [72] NeurIPS’24 $\checkmark$ $\checkmark$ 31.3 37.1ms 76.1 91.9 61.1 91.8 91.3 97.2 84.9 69.5 60.5
E2E-OSDet - $\checkmark$ $\checkmark$ 27.5 20.9ms 77.7 90.7 64.3 91.8 92.1 97.8 85.7 70.3 61.4
Table 3: Ablation analysis of E2E-OSDet.
[…]
(a) 83.4 66.4 58.0
(b) $\checkmark$ 84.5 67.1 58.9
(c) $\checkmark$ 84.9 68.6 59.6
(d) $\checkmark$ 84.6 68.9 60.0
(e) $\checkmark$ $\checkmark$ 85.5 69.9 61.0
(f) $\checkmark$ $\checkmark$ 84.9 69.5 60.7
(g) $\checkmark$ $\checkmark$ 85.2 70.0 60.9
(h) $\checkmark$ $\checkmark$ $\checkmark$ 85.7 70.3 61.4
Ablation Experiment. As shown in Table 3(a), we adopt CSSA [2] as our baseline model, which serves as the foundation for independently evaluating the contribution of each proposed module. The results (see Table 3 (a), (b), and (c)) indicate that all three modules lead to performance gains, among which the AFM module provides the most substantial improvement in detection accuracy. We further assess the interoperability of the modules. As shown in Table 3 (e), (f), and (g), the modules are compatible with one another, and their integration does not result in any performance degradation. Finally, we integrate all modules (see Table 3 (h)), resulting in additional gains in overall model performance.
Figure 6: (a) Effect of different handcrafted features. (b) Effect of different scanning mechanisms. (c) Effect of varying ratios of handcrafted features. (d) Effect of different attention regions. (e) Performance of E2E-OSDet across various detectors. (f) Training time of the single-source method (using SAR images). (g) Training time of the multi-source method (using optical and SAR images).
Figure 7: Visualization results of handcrafted features. SSIM represents structural similarity.
Impact of Filter Augment and Scanning Mechanisms. To assess the effectiveness of the filter-augmented inputs, we tested a range of conventional feature descriptors. As shown in Figure 6 (a) and Figure 7, the results indicate that incorporating handcrafted features significantly enhances detection performance. This finding suggests that mapping pixels into a handcrafted feature space effectively reduces the distribution gap between optical and SAR data. Notably, the Grad feature yields superior performance due to its ability to narrow domain differences and capture multi-scale information. We further evaluate the impact of the CMIM module under different scanning mechanisms [15]. As shown in Figure 6 (b), CMIM consistently enhances performance across all scanning strategies. In addition, we analyze the effects of the proportion of handcrafted feature input and the use of area attention. As shown in Figure 6(c) and (d), these hyperparameters exhibit marginal effects on performance and do not substantially influence the final results.
4.3 Qualitative Evaluation
We present visual comparisons of detection results from various multi-source methods. As shown in Figure 8, the proposed E2E-OSDet consistently outperforms other methods in complex scenes. In particular, in regions with limited optical information, such as cloud-covered, snow-covered, low-light, or low-resolution areas, the incorporation of SAR data significantly enhances target localization accuracy. These experimental results demonstrate the proposed dataset’s effectiveness in evaluating the robustness and adaptability of detection models across diverse scenarios. Additional qualitative results, including detection outputs and heatmaps, are provided in Appendix D.
5 Discussion
Summary. This work addresses the lack of publicly available, high-quality datasets for optical–SAR fusion object detection by introducing the M4-SAR dataset, the first large-scale dataset specifically tailored for this task. Furthermore, we propose E2E-OSDet, a novel detection framework that demonstrates strong robustness across diverse and challenging scenarios.
Limitations Analysis and Future Work. Despite the significant advantages of M4-SAR in terms of scale, diversity, and scene coverage, there remain several limitations that merit further exploration: First, a temporal gap exists between the acquisition of optical and SAR images, particularly affecting a minority of dynamic targets by introducing spatial inconsistencies. Future work could explore robust detection methods to handle temporal misalignment [68; 3; 10]. Second, although a semi-supervised annotation strategy is employed, SAR labels still rely on guidance from optical images, which can be unreliable under occlusion or in the presence of blurred boundaries. To address this, future research could introduce strategies for modeling missing modalities [57; 34] to enhance the model’s adaptability to weak or absent inputs. Third, the current dataset focuses on six representative static object categories and lacks more dynamic or fine-grained object categories. Expanding the object taxonomy and annotation granularity would benefit a wider range of remote sensing applications [58; 47]. Finally, although E2E-OSDet offers a strong performance baseline, its computational complexity hinders its deployment on resource-constrained edge devices. Future efforts could focus on developing lightweight multi-source detection frameworks (e.g., DecoupleNet [37], SAMNet [30]) to maintain high accuracy while improving inference efficiency.
Usefulness of M4-SAR. In benchmark evaluations, our proposed M4-SAR has demonstrated its effectiveness in leveraging the complementary characteristics of optical and SAR images to improve object detection accuracy. We believe that this dataset is not only suitable for multi-source fusion object detection but is also broadly applicable to cross-source reconstruction, multi-source alignment, multi-source image dehazing, and other related tasks.
Figure 8: Qualitative comparison between the proposed E2E-OSDet and six fusion detection methods.
6 Conclusion
This paper presents the M4-SAR dataset and the E2E-OSDet method to tackle two persistent challenges in optical–SAR fusion object detection: the scarcity of high-quality, large-scale datasets and the absence of a unified and effective algorithmic framework. We also develop MSRODet, a comprehensive open-source toolkit for multi-source rotated object detection that integrates mainstream fusion detection algorithms and enables fair performance comparisons on the M4-SAR dataset. We expect that M4-SAR, as a benchmark dataset, will substantially advance research in multi-source remote sensing detection, while E2E-OSDet, as the first fusion framework specifically designed for optical–SAR scenarios, will serve as a strong baseline for future studies.
Appendix
This appendix is organized as follows:
  • Appendix A provides comprehensive details on the structure and annotation protocol of the proposed M4-SAR dataset.
  • Appendix B offers additional background for this study, including a review of existing SAR datasets, SAR-based object detection methods, and multi-source fusion object detection approaches.
  • Appendix C presents an in-depth description of the E2E-OSDet architecture, including the Filter Augment Module (FAM), Cross-modal Mamba Interaction Module (CMIM), and Area-attention Fusion Module (AFM), emphasizing their roles in enhancing cross-modal feature fusion.
  • Appendix D outlines implementation details, ablation studies, and visualization results, further validating the effectiveness and robustness of the proposed method under diverse conditions.
Appendix A M4-SAR
A.1 Dataset and Label Format
We introduce M4-SAR, a novel multi-source remote sensing dataset designed to advance research in multi-source object detection. M4-SAR combines high-resolution optical and synthetic aperture radar (SAR) images, capturing complementary characteristics to address challenges posed by varying illumination, adverse weather conditions, and modality differences. The dataset is carefully organized into three subsets: training, validation, and testing, each comprising paired optical–SAR images with spatially aligned and consistently annotated ground truths.
Figure 9: An example of annotation in our M4-SAR dataset.
Annotations are provided in a flexible .txt format, where each object instance is represented by a quadrilateral defined by four vertices: $[category,(x_{1},y_{1}),(x_{2},y_{2}),(x_{3},y_{3}),(x_{4},y_{4})]$ (see Figure 9). All coordinates are normalized to the range $[0,1]$ relative to the image dimensions, allowing precise localization of arbitrarily oriented objects—an essential feature for detecting rotated or skewed targets common in remote sensing images. The dataset includes six semantically diverse categories: bridge, harbor, oil tank, playground, airport, and wind turbine. These categories cover critical infrastructure and industrial facilities, representing real-world scenes across diverse geographic and environmental conditions. M4-SAR is designed to drive progress in multi-source object detection by offering a challenging and comprehensive benchmark. Through the integration of optical and SAR modalities with orientation-aware, high-precision annotations, M4-SAR addresses key limitations of existing datasets and provides a solid foundation for developing robust, generalizable models for remote sensing applications. More sample visualization of the M4-SAR dataset is shown in Figure 10.
A.2 Societal Impact and Ethics Consideration
First, the M4-SAR dataset offers a new perspective on multi-source fusion detection, which can bring about positive societal effects in various applications such as disaster monitoring and national defense. A typical negative societal impact during data collection involves the inclusion of sensitive areas, such as military zones or private properties. We have carefully reviewed this aspect, and we ensured that our acquisition area was limited to non-military and non-private areas.
Figure 10: Sample visualization of the M4-SAR dataset.
A.3 License Declaration
The proposed M4-SAR dataset and E2E-OSDet are freely available under the CC BY 4.0 license agreement: https://creativecommons.org/licenses/by/4.0/.
Appendix B Related Work
B.1 Datasets
The development of SAR datasets for object detection faces significant challenges, primarily due to the high cost of data acquisition and the complexity of fine-grained annotation. As a result, existing SAR datasets, such as HRSID [61], SADD [73], SSDD [75], and SIVED [29] are typically limited to a few easily recognizable categories (e.g., ships) and contain fewer than 10,000 images with minimal class diversity (see Table 1). These limitations hinder the development of robust and generalizable models for remote sensing applications.
Recent efforts have sought to overcome these constraints. Wang et al. [59] introduced the Ship-Dataset, comprising 39,729 ship image patches from 102 Gaofen-3 and 108 Sentinel-1 scenes. While large in scale, this dataset lacks semantic diversity. Xia et al. [63] proposed the MSAR dataset, expanding the category set to include airplanes, tanks, bridges, and ships. To address data scarcity at scale, Li et al. [28] released SARDet-100K, containing 116,598 images and 245,653 labeled instances spanning six categories: bridges, harbor, oil tanks, playground, airport, and wind turbines. More recently, Wu et al. [62] introduced FAIR-CSAR with fine-grained annotations and coverage across multiple domains. Zhang et al. [76] further enhanced annotation efficiency in RSAR by leveraging weakly supervised strategies based on SARDet-100K, incorporating rotated bounding boxes.
While these developments mark significant progress, from early small-scale, single-category datasets to large-scale, multi-class benchmarks, most existing SAR datasets remain single-source. This limits their capacity to capture complementary information from different modalities, which is crucial for robust object detection in real-world remote sensing scenarios. To address this gap, Wang et al. [55] proposed OGSOD-1.0, the first dataset to align optical and SAR images. However, its evaluation protocol relies solely on SAR data, restricting its applicability to comprehensive optical–SAR fusion.
Therefore, there is a critical need for a large-scale, high-quality multi-source dataset that fully supports end-to-end optical–SAR fusion. Such a dataset would enable robust and generalizable multi-modal detection under diverse environmental conditions, significantly advancing the state of remote sensing perception.
B.2 SAR Object Detection
Synthetic aperture radar (SAR) images are inherently degraded by multiplicative speckle noise and structural artifacts, which complicate the extraction of discriminative features for object detection. Early methods addressed these challenges using handcrafted feature descriptors, including the Canny Edge Detector [1], Wavelet Scattering Transform (WST) [40], Haar-like features [52], Histogram of Oriented Gradients (HOG) [5], and Gradient by Ratio Edge (Grad) [24].
Canny Edge Detector (Canny): A classical algorithm that combines Gaussian smoothing, gradient estimation, and non-maximum suppression to extract robust edges while suppressing noise. Its utility in SAR edge detection has been validated in previous studies [60; 11].
Wavelet Scattering Transform (WST): WST captures hierarchical, multi-scale features by preserving both high- and low-frequency components, enabling robust representations of small objects in noisy SAR environments.
Haar-like Features (Haar): Originally used for face detection, Haar-like features employ predefined templates to capture edges and corners, and have been adapted for SAR target recognition due to their computational efficiency and robustness to noise.
Histogram of Oriented Gradients (HOG): HOG encodes local shape and structural patterns through gradient orientation histograms, offering robustness to speckle noise and invariance to illumination changes in SAR images.
Gradient by Ratio Edge (Grad): Specifically designed for multiplicative noise, Grad leverages a Ratio-of-Averages operator [51] to extract edges, and has been integrated into SAR-specific descriptors such as SAR-HOG [6] and SAR-SIFT [46] for improved performance.
More recently, deep learning has transformed object detection, with architectures such as R-CNN [13; 16], YOLO [21; 22; 56; 53], and DETR demonstrating strong generalization in natural images. However, SAR-based object detection remains particularly challenging due to speckle noise and limited semantic cues, which degrade feature discriminability. To address this, recent approaches have explored SAR-specific architectures. For instance, the SFS-Conv module [23] incorporates diversion sensing and fractional Gabor transforms to enhance frequency-domain representation, while SAR-JEPA [24] introduces a joint embedding framework that predicts multi-scale gradient features via masked local reconstruction.
Despite these advances, the majority of existing methods are designed for single-source SAR data and fail to exploit the complementary strengths of auxiliary modalities such as optical images. This limitation highlights the need for multi-source fusion frameworks capable of leveraging optical–SAR synergy to mitigate noise, enrich semantic context, and improve detection robustness in complex remote sensing scenarios.
B.3 Multi-source Fusion Object Detection
Deep neural networks, owing to their strong nonlinear modeling capacity, have significantly advanced multi-source fusion object detection [45; 2; 44]. Recent studies have tackled key challenges in cross-modal integration. For example, Yuan et al. [67] and Fang et al. [42] proposed alignment strategies to mitigate spatial misalignments between heterogeneous modalities. He et al. [18] investigated semantic conflicts and complementarity across sources to enhance fusion effectiveness. Zeng et al. [69] introduced a contour enhancement module to extract modality-specific spectral features from visible and infrared images, emphasizing spatial diversity and fine-grained structures. Despite these advances, many existing frameworks suffer from complex training procedures, limiting scalability and real-world deployment. To address this, Zhang et al. [72] proposed E2E-MFD, an end-to-end architecture that streamlines the fusion pipeline through a single-stage training process while achieving strong performance.
However, the majority of prior research has focused on visible–infrared or multispectral-RGB fusion, while optical–SAR fusion remains relatively underexplored. This setting poses unique challenges due to significant disparities in feature distributions and sensing mechanisms. Optical images offer rich color and texture cues but are sensitive to lighting and weather, whereas SAR data provides geometric and structural information resilient to environmental changes. Effectively integrating these complementary properties is essential for robust object detection in complex remote sensing scenarios. The lack of large-scale, high-quality optical–SAR datasets and specialized fusion algorithms highlights an urgent need for new frameworks tailored to this modality pair. Bridging this gap is critical for enabling generalizable, noise-resilient, and efficient multi-source object detection in real-world applications.
Figure 11: Mechanisms for scanning in different directions.
Appendix C More details of E2E-OSDet
C.1 Filter Augment Module (FAM)
To address the inherent domain disparities between optical and SAR images, caused by differences in imaging mechanisms and feature distributions, we propose the Filter Augment Module (FAM), a novel preprocessing framework designed to harmonize multi-source inputs for robust object detection. FAM incorporates a suite of classical image filtering operators, including Histogram of Oriented Gradients (HOG) [5], Canny Edge Detector [1], Haar-like Features [52], Gradient by Ratio Edge (Grad) [24], and Wavelet Scattering Transform (WST) [40]. These filters extract complementary low-level cues, such as edges, gradients, and textures, under diverse imaging conditions. By projecting the sparse, low-dimensional features of both optical and SAR images into a high-dimensional, discriminative feature space, FAM reduces cross-modal domain gaps and enhances the alignment of modality-specific characteristics. Specifically, given an optical image $X$ and a SAR image $Y$, FAM generates a set of enhanced representations $\{X_{i}\}$ and $\{Y_{i}\}$ for each filter $i\in\{\text{WST},\text{Canny},\text{Haar},\text{HOG},\text{Grad}\}$, defined as:
$X_{i}=X+\alpha*\mathbb{FAM}_{i}(X),\quad i\in\{WST,Canny,Haar,Hog,Grad\},$ (1)
$Y_{i}=Y+\alpha*\mathbb{FAM}_{i}(Y),\quad i\in\{WST,Canny,Haar,Hog,Grad\},$ (2)
where $\mathbb{FAM}_{i}(\cdot)$ denotes the $i$-th filter-specific augmentation function, and $\alpha$ represents a set of learnable scalars that balance the contributions of the original and filtered features for the optical and SAR modalities, respectively. This dual-branch enhancement scheme preserves the structural and textural integrity of each modality while injecting noise-robust, multi-scale edge and shape cues derived from classical filters. By aligning feature representations across modalities at the preprocessing stage, FAM facilitates more effective optical–SAR fusion. It addresses the limitations of conventional preprocessing pipelines and significantly enhances the generalization capability of multi-source object detection models under diverse and challenging remote sensing conditions.
C.2 Cross-modal Mamba Interaction Module (CMIM)
Existing multi-source fusion approaches primarily emphasize spatial-level feature integration, often neglecting the substantial domain discrepancies between heterogeneous modalities such as optical and SAR images. These domain shifts pose a significant challenge to effective cross-modal fusion. The Mamba framework [14], with its state-space modeling and efficient sequence processing capabilities, has demonstrated strong potential for capturing long-range dependencies. This makes it well-suited for mitigating inter-modal inconsistencies. However, existing Mamba-based fusion methods [70; 8] typically involve intricate architectures and computationally intensive interaction designs, which limit their scalability and practical deployment. To overcome these limitations, we propose the Cross-modal Mamba Interaction Module (CMIM), a lightweight yet powerful module tailored to align optical and SAR features by leveraging Mamba’s long-sequence modeling strength. Given paired image features $X_{i}$ and $Y_{i}$, we first flatten them into one-dimensional sequences in row-major order:
$X_{i}=[a_{1},a_{2},a_{3},\dots,a_{n-1},a_{n}],\quad Y_{i}=[b_{1},b_{2},b_{3},\dots,b_{n-1},b_{n}].$ (3)
Traditional methods concatenate these sequences serially:
$Z_{\text{concat}}=[a_{1},a_{2},a_{3},\dots,a_{n-1},a_{n},b_{1},b_{2},b_{3},\dots,b_{n-1},b_{n}].$ (4)
However, these sequence modeling approaches often fail to establish explicit local correspondences between modalities, leading to suboptimal cross-modal alignment and feature fusion.
To address this issue, we introduce an Interleaved Input Rearrangement (IIR) mechanism that enforces fine-grained spatial correspondence by interleaving paired features in a one-to-one fashion. Specifically, given vectorized feature sequences $X_{i}\in\mathbb{R}^{L\times d}$ and $Y_{i}\in\mathbb{R}^{L\times d}$ from the optical and SAR modalities, respectively, IIR constructs a unified sequence $Z_{i}\in\mathbb{R}^{2L\times d}$ by interleaving the features as:
$Z_{\text{concat}}=[a_{1},b_{1},a_{2},b_{2},a_{3},b_{3},\dots,a_{n-1},b_{n-1},a_{n},b_{n}].$ (5)
This interleaving ensures that corresponding positional features from both modalities are adjacent, enabling Mamba’s state-space modeling to directly capture semantic and structural alignments, thus enhancing domain coherence. The resulting interleaved sequence is subsequently fed into a directional scanning mechanism (see Figure 11 [19]), which applies Mamba blocks along both horizontal and vertical directions. This bi-directional sequence modeling allows CMIM to capture long-range dependencies while preserving fine-grained local correspondence, facilitating deeper cross-modal interaction. Finally, the output sequence is reshaped back into the 2D spatial domain and fused with residual skip connections to retain original modality-specific information. Through this design, CMIM effectively bridges modality gaps and enhances the robustness of multi-source object detection under challenging remote sensing conditions. :
$Z_{\text{scan}}=\mathbb{SCAN}_{i}(Z_{\text{concat}})\quad i\in\{Bid,Zorder,Zigzag,Hilbert\}.$ (6)
The CMIM then transforms the scanned sequence into enhanced feature representations:
$(F^{opt}_{j},F^{sar}_{j})=\mathbb{CMIM}(Z_{\text{scan}}),\quad j\in\{3,4,5\}.$ (7)
By positioning optical and SAR features adjacent along the temporal axis, the Interleaved Input Rearrangement (IIR) mechanism enables synchronous semantic enhancement within short time steps, effectively mitigating the risk of feature divergence during training. This temporal alignment facilitates efficient parameter sharing across modalities, stabilizes the training process, and promotes consistent gradient flow. As a result, IIR significantly enhances cross-modal feature alignment and consistently outperforms conventional fusion strategies in complex remote sensing scenarios.
C.3 Area-attention Fusion Module
To enhance the discriminative capacity of fused optical–SAR features across spatially diverse regions, we propose the Area-attention Fusion Module (AFM), a novel mechanism that emphasizes salient local structures and amplifies responses in critical target regions. AFM exploits the complementary characteristics of optical and SAR modalities, enabling robust and adaptive feature integration for multi-source object detection. Inspired by the efficient spatial modeling techniques in [48; 12], AFM partitions the input feature map of resolution $(H,W)$ into $n$ non-overlapping blocks of size $(H/n,W)$ or $(H,W/n)$, without requiring explicit window segmentation. This lightweight partitioning is implemented via a simple reshaping operation, preserving spatial continuity while substantially improving computational efficiency. The fused feature representation is computed as:
$F^{fus}_{j}=\mathbb{AFM}(F^{opt}_{j},F^{sar}_{j}),\quad j\in\{3,4,5\},$ (8)
where $F^{opt}_{j}$ and $F^{sar}_{j}$ denote the optical and SAR features (e.g., outputs from the Cross-modal Mamba Interaction Module, CMIM), and $\mathbb{AFM}()$ denotes the attention-driven fusion function that selectively emphasizes salient regions. By adaptively weighting localized features, AFM improves the model’s capability to detect targets in complex remote sensing scenes characterized by varying object scales and cluttered backgrounds. As illustrated in Figure 5, AFM achieves efficient and effective cross-modal fusion, surpassing conventional spatial fusion strategies that often overlook regional saliency and incur higher computational overhead.
C.4 Loss Function
In the proposed E2E-OSDet framework, we employ a composite loss function comprising Binary Cross-Entropy (BCE) and Probabilistic IoU (ProbIoU) [32] to jointly optimize classification and localization. The overall loss is defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{reg}}+\mathcal{L}_{\text{dfl}}+\mathcal{L}_{\text{cls}},$ (9)
where $\mathcal{L}_{\text{reg}}$ denotes the bounding box regression loss, $\mathcal{L}_{\text{dfl}}$ represents the distribution focus loss, and $\mathcal{L}_{\text{cls}}$ corresponds to the classification loss. This formulation balances localization accuracy and classification confidence, ensuring end-to-end optimization of the optical–SAR fusion detection pipeline.
Figure 12: Visualization of handcrafted features on SAR images, with SSIM denoting structural similarity. Features are average-pooled and represented as a single channel for visualization.
Figure 13: Comparison of detection results across different methods (zoom in for detail).
Appendix D Experimental Results
D.1 Implementaion Details
To facilitate robust evaluation and accelerate progress in multi-source object detection, we introduce MSRODet, a comprehensive and extensible toolbox built upon the Ultralytics framework [21]. MSRODet integrates a wide range of fusion-based detectors, including CFT [42], CLANet [18], CSSA [2], CMADet [45], ICAFusion [44], MMIDet [69], and our proposed E2E-OSDet, alongside state-of-the-art single-source baselines such as YOLOv5 [21], YOLOv6 [22], YOLOv8 [21], YOLOv9 [56], YOLOv10 [53], and YOLOv11 [21]. To ensure fair comparisons, all multi-source fusion methods adopt YOLOv11-S as the unified backbone and neck architecture. We also extend these methods to other YOLO variants to evaluate their framework compatibility while maintaining consistent model capacity.
For oriented object detection, crucial in remote sensing applications, we replace the standard detection heads with the oriented bounding box (OBB) head from YOLOv8 [21], enabling accurate localization of arbitrarily rotated targets. Training is performed using a momentum-based SGD optimizer (initial learning rate: 0.01, momentum: 0.937, weight decay: 5e-4) for 300 epochs with a batch size of 64. All input images are resized to 512$\times$512, and experiments are conducted on two NVIDIA RTX 4090 GPUs with mixed-precision (FP16) training. To eliminate pretraining bias, all fusion models are trained from scratch.
For benchmarking, we also evaluate single-source baselines implemented via the MMRotate toolkit [41], including Rotated FCOS [49], Rotated ATSS [74], Oriented R-CNN [64], Oriented RepPoints [25], RTMDet [38], PSC [66], and LSKNet [27]. All methods use a ResNet-50 backbone [17] pre-trained on ImageNet, and are optimized using AdamW [33] (learning rate: 2e-4, weight decay: 0.05) for 72 epochs with a batch size of 32 and the same 512$\times$512 input resolution. This unified and standardized setup ensures fair, reproducible, and extensible comparisons across a wide spectrum of single-source and multi-source object detection methods, establishing MSRODet as a robust benchmark platform for optical–SAR fusion in remote sensing.
D.2 Visualization Results for Handcrafted Feature
To assess the effectiveness of handcrafted features in mitigating domain disparities between optical and SAR images, we examine a set of classical feature descriptors, including the Canny Edge Detector [1], Wavelet Scattering Transform (WST) [40], Haar-like Features [52], Histogram of Oriented Gradients (HOG) [5], and Gradient by Ratio Edge (Grad) [24]. We employ the Structural Similarity Index Measure (SSIM) to quantitatively evaluate the structural alignment between optical and SAR images in the handcrafted feature space. As shown in Figure 12, these handcrafted descriptors significantly enhance cross-modal structural similarity compared to the original pixel domain, effectively bridging the intrinsic imaging differences between modalities.
Notably, the Grad feature space [24] achieves the highest SSIM scores, significantly outperforming other descriptors by capturing noise-resilient edge information that is particularly critical for SAR images. This result highlights Grad’s superior capacity for aligning multi-source features, especially under challenging conditions characterized by multiplicative speckle noise in SAR data. Motivated by this observation, our method prioritizes the integration of Grad-based handcrafted features within the Filter Augment Module (FAM) to enhance the structural alignment and fusion of optical–SAR representations, thereby improving the robustness and discriminability of multi-source object detection in complex remote sensing environments.
D.3 Detection Result Visualizations
To elucidate the effectiveness of our multi-source fusion framework for optical–SAR object detection, we present a comprehensive visual comparison of detection outcomes across diverse input modalities and fusion strategies. As shown in Figure 13, our method effectively harnesses the complementary characteristics of optical (rich texture and color) and SAR (robust structural information) images, yielding sharper target boundaries and higher detection confidence relative to existing fusion baselines. In particularly challenging scenarios—including low-contrast environments, cluttered urban scenes, and adverse weather conditions—our approach exhibits strong resilience, significantly improving both recognition accuracy and localization precision. These visual comparisons highlight the critical role of cross-modal alignment and deep semantic integration, underscoring the practical applicability and generalization capability of our method in real-world remote sensing tasks.
D.4 Heatmap Visualizations
To assess the effectiveness of our model in leveraging multi-source information, we present attention heatmap visualizations comparing the response distributions of various methods on representative samples from our optical–SAR dataset. As illustrated in Figure 14, our proposed E2E-OSDet produces highly concentrated and spatially aligned attention maps, accurately focusing on target regions while suppressing irrelevant background activations. This indicates a strong ability to extract semantically meaningful and spatially coherent features from heterogeneous modalities. In contrast, baseline fusion methods yield diffuse or misaligned attention responses, often exhibiting attention drift or overlooking low-contrast objects, especially in cluttered or noisy scenes. These qualitative results emphasize the importance of cross-modal alignment and semantic integration, and demonstrate that E2E-OSDet significantly improves attention precision, detection robustness, and generalization to complex remote sensing environments.
Figure 14: Grad-CAM [43] heatmaps of the proposed E2E-OSDet and six fusion detection methods.